The cost of 8 kg of potatoes and 10.kg of onions is Rs.46.The cost of 10.kg of potatoes and 5 kg of onions is Rs.35.Find the price of potato per kg and onion per kg?.Find the solution of equation??
I took 8x+10y=46
10x+5y=35
Then what should do?



Answer by advanced_Learner(501)

8x+10y=46
10x+5y=35
8x+10y=46
-20x-10y=-70
-12x=-24
x=2 rbs
8x+10y=46
16+10y=46
y=3 rbs